# MetaChest: Generalized few-shot learning of patologies from chest X-rays

### v1

@@ 1 Introduction
In recent decades, deep learning has revolutionized medical image analysis, particularly in the field of radiology [35, 24, 8]. Deep neural networks have enabled the processing of large volumes of radiological data, the extraction of complex features, and the development of models that can enhance the accuracy of medical diagnoses. Despite these advances, a major challenge arises when only limited annotated data are available, as deep learning models typically require large amounts of labeled data to achieve strong performance. This issue is especially relevant to tasks such as pathology classification in chest X-rays, where labeled data can be scarce and difficult to obtain. To address this limitation, early research has explored the use of the standard few-shot classification (SFSC) paradigm, which aims to train models capable of generalizing to new classes using only a few labeled examples per class. However, this paradigm differs significantly from the way pathologies manifest in practice. From a clinical perspective, the objective is not merely to classify entirely new disease categories, but rather to distinguish between a combination of known pathologies and previously unseen ones. This highlights the need for approaches that are more closely aligned with the complexities of clinical settings.
@@ 2.1 Deep learning for chest X-ray classification
Deep neural networks, coupled with large-scale datasets, have enabled significant progress in several computer vision fields. Over the past few years, in an effort to take advantage of deep neural networks, datasets of increasingly specialized domains have been made publicly available. For instance, in the medical domain, multiple chest X-ray datasets have been introduced, such as CheXpert [19], Chest X-ray8 [57], Chest X-ray14 [57], MIMIC [21], MIMIC-CXR-JPG [22], OpenI [11], and PadChest [5]. These datasets have been fundamental in the development of deep learning models for chest X-ray analysis and generation tasks. In contrast to ImageNet [12], the scale of these datasets is at least one order of magnitude smaller. In addition, the distributions of these datasets are greatly heterogeneous, e.g. the number and kind of pathologies, the class imbalance, the collection and labeling procedure, the quality of the images, and the patient population.
With the introduction of these datasets, several studies addressing pathology classification from chest X-rays using deep learning started to appear. For binary classification (i.e. presence or absence), Lakhani & Sundaram [28] focused on tuberculosis identification, whereas Mabrouk et al. [31] targeted pneumonia identification. Since the X-ray of a patient may exhibit signs of multiple diseases, the identification of pathologies from chest X-rays has often been formulated as a multi-label classification problem. For instance, Baltruschat et al. [3] used a ResNet-50 [15] architecture to classify fourteen pathologies in the ChestX-ray14 dataset, where each X-ray could be assigned to more than one pathology. Similarly, Irvin et al. [19] compared various ConvNet architectures for multi-label chest X-ray classification using CheXpert, finding that DenseNet121 outperformed ResNet152 [15], Inception-v4 [49], and SE-ResNeXt101 [17].
@@ 2.2 Standard transfer learning
Transfer learning (TL) is a cornerstone of deep learning for image analysis, since it can reduce the amount of data and the computational resources required to train a model for a target task by leveraging the representations learned from one or multiple source tasks. In TL, multiple strategies to adapt the representations from the source task to the target task have been proposed. In practice, the most widely used transfer strategy has been standard transfer learning, which consists of pre-training models using a conventional batch-based training (as opposed to other training schemes). Specifically, ImageNet pre-training has been a standard practice for a wide variety of natural image tasks, including classification [50, 26], segmentation [14, 41] and object detection [29].
@@
Due to the widespread use of ImageNet in practice, multiple works have studied the transferability of the learned representations from ImageNet to other natural image tasks [40, 61, 27, 63]. Surprisingly, even though some studies have suggested that the source and target datasets must be closely related for an effective knowledge transfer [61], ImageNet pre-training has been used with seeming success for wildly dissimilar image domains (e.g. medical images [57, 38, 24]). In contrast, there are transferability studies in specific domains where ImageNet pre-training has not provided any improvement over random initialization [38, 35, 8].
@@
Other transferability studies have focused on analyzing the effect of the architecture size and the scale of the training dataset on the effectiveness of STL. In intra-domain scenarios, where the source and target datasets are closely related, studies have been mainly focused on natural image datasets. For example, Kolesnikov et al. [26] and Zhai et al. [62] analyzed how the pre-training dataset size and the architecture depth influence knowledge transfer when both the source and target datasets are composed of natural images. The results from these studies have consistently shown better performance with larger architectures and pre-training datasets.
@@
In inter-domain scenarios, where the source and target datasets belong to different domains (e.g., natural images and chest X-rays), existing studies are scarce, not very systematic and report mixed results. Raghu et al. [38] did not find significant differences on chest X-ray and retinal image classification performance using a ResNet-50 architecture when comparing ImageNet-1k pre-training with random initialization. Ke et al. [24] studied the effect of ImageNet-1k pre-training on chest X-ray classification performance using ConvNet architectures of different sizes. Their results showed a slight performance improvement when using deeper pre-trained architectures. On the other hand, Mustafa et al. [35] studied the influence of ImageNet-1k, ImageNet-21k, and JFT-300M pre-training on classification performance using ResNets of different sizes. The target tasks considered in this study were cancer identification from mammograms, pathology classification from chest X-rays, and skin condition from dermatological images. The results were far from conclusive, observing performance improvements with larger pre-training datasets and architectures only in some target tasks. Similarly, Cherti & Jitsev [8] carried out a comparative study of ResNet models pre-trained on ImageNet-1k, ImageNet-21k, and a combination of different chest X-ray datasets for pathology classification. They reported small improvements in performance when models were pre-trained on larger source datasets and transferred to larger target datasets. However, no performance improvement was observed when transferring for smaller target datasets, regardless of the size of the pre-training dataset.
@@ 2.3 Few-shot classification
Meta-learning is a transfer learning strategy which aims to generate models that can be quickly adapted to new tasks [52]. As opposed to STL, in meta-learning, new tasks are commonly known as episodes and are typically small with respect to both the number of classes and the number of examples per class. The most widely studied problem in meta-learning is few-shot classification, which is a multi-class classification problem where a few examples per class are available for training (typically, 1 or 5).
The earliest works in meta-learning proposed methods for multi-class classification on natural image datasets [13, 55, 46, 48, 1], such as MiniImageNet [55] and FC100 [47]. These datasets are reduced versions of ImageNet and CIFAR100 and were created to facilitate episodic training.
More recent works have applied meta-learning to domain-specific problems, particularly through datasets of various medical imaging modalities. For instance, meta-learning methods have been studied in skin disease classification from dermatological images [32], COVID-19 classification from chest CT scans [7], and cancer classification from histological images [33, 45]. Moreover, meta-learning methods have also been used for image segmentation in CT scans, magnetic resonance images [51], and dermatological images [25].
@@ 3 MetaChest dataset
Clinical data collection is a complex process involving several tasks which can require a considerable amount of time and resources. Data labeling is one of the tasks that can generate greater variability among chest X-ray datasets; the rightmost column in Table 1 summarizes the labeling strategy employed by each dataset. Most datasets derived annotations automatically from radiology reports using natural language processing (NLP) methods, with the exception of PadChest which was annotated by expert radiologists. The specific strategy and tool used for annotating the chest X-rays directly influences the distribution of labels. For instance, MIMIC provides two sets of labels that have different distributions: one generated by NegBio [37] and another by CheXpert [19].
@@
Table 1: Comparison of publicly available chest X-ray datasets. Here, MIMIC [21] refers to MIMIC-CXR-JPG [22].
@@ 3.1 Data
The frequency of each pathology in MetaChest is shown in Figure 1. As can be observed, there is a pronounced class imbalance, with the most frequent pathology (Effusion) occurring nearly two orders of magnitude more often than the least frequent one (Hernia). With respect to labeling, MetaChest has a label cardinality (average number of labels per image) of 1.84 and a label density (label over the total number of labels; see Tsoumakas et al. [53]) of 0.12.
@@ 3.2 Meta-learning partition
Although Cherti & Jitsev [8] used a chest X-ray dataset that combined multiple datasets, it is not publicly available, its generation procedure is not described, and it does not provide appropriate partitions for meta-learning evaluation. Similarly, TorchXRayVision [10] is a library that allows the combination of different chest X-ray datasets, but it does not consider SFSL scenarios either. Conversely, MetaChest employs a disjoint class partition, enabling experimentation in SFSL settings. In addition, $\mathcal{C}_{meta-tst}$ is composed of the classes with the fewest examples available across the four original datasets, which is convenient for evaluating classification methods on images collected from multiple hospitals.
@@ 4.1 Standard transfer learning
Standard transfer learning (STL) is the most widely studied and spread strategy for computer vision tasks. When performing STL, we can identify two main stages [20]:
  • 1.
@@ 4.2 Meta-learning
Meta-learning (MTL) is an alternative paradigm to STL, which aims to obtain models that can adapt to novel tasks with unseen classes and very few examples per class [20]. In other words, the objective of MTL is to achieve a more efficient transfer in terms of data. Similar to STL, the transfer process of MTL consists of a pre-training stage followed by an adaptation stage. Although in MTL these stages are commonly referred to as meta-training and meta-test [39], for the sake of consistency and clarity, we use the terms pre-training and adaptation for both STL and MTL. In this work, we are interested in two MTL formulations for classification: standard few-shot classification (SFSC) and generalized few-shot classification.
@@ 5.2.1 ProtoNet-ML
ProtoNet [46] is a multi-class classification method that has been widely studied in the SFSC literature. In this work, we propose an extension to handle multi-label classification, which we call ProtoNet-ML. Following the original method, ProtoNet-ML computes a $D$-dimensional prototype $\mathbf{z}_{c}\in\mathbb{R}^{D}$ for each class $c$ as follows:
@@
Figure 4: Representation subspaces for multi-class ProtoNet (left) and ProtoNet-ML (right). In multi-class ProtoNet, a representation vector $g_{\phi}(\mathbf{x})$ is associated with only one prototype, whereas in ProtoNet-ML, a representation vector $f_{\phi}(\mathbf{x})$ may be associated with one or more prototypes.
@@ 5.2.2 BatchBased
BatchBased is a method inspired by [6] that employs STL-based training while maintaining MTL-based episode evaluation. On top of the backbone $f_{\phi}$, BatchBased adds a head module $g_{\varphi}$ (single fully connected layer) with trainable parameters $\varphi$. The probability distribution for a test example $\mathbf{x}$ given a class $c$ is computed as:
@@ 6 Results and discussion
In this section, we analyze the adaptation process of different models across various formulations of the multi-label classification task using few examples per pathology. First, we compare how two distinct learning paradigms leverage ImageNet pre-training. Next, we examine their behavior across a broad set of few-shot learning tasks designed to reflect challenges commonly encountered in medical settings. Finally, we investigate factors that influence the adaptation process, including image resolution, and variations in architectural connectivity patterns.
@@ 6.1 Experimental setup
Parameter Configuration Data     Distribution Complete     Image size 384 Task     Training batch size 64     Training $n$-way, $k_{trn}$ $k_{tst}$ 3, 30, 30     Validation $n$-way, $k_{trn}$ $k_{tst}$ 3, 30, 30     Test $n$-way, $k_{trn}$ $k_{tst}$ 3, 30, 30 Backbone     Architecture MobileNetV3Small075     pre-training I1K Training     Meta-trn, meta-val, meta-tst episodes 1,000, 100, 10,000     Max epochs 150     Optimizer AdamW     Stop metric, patience $HM$, 10     Float precision 16bit BatchBased     Meta-trn LR 0.0001     Meta-val $t_{steps}$, $ptc_{trn}$, $lr_{head}$ 100, 0.5, 0.05     Meta-tst $t_{steps}$, $ptc_{trn}$, $lr_{head}$ 100, 0.5, 0.05 ProtoNet     Encoder type, size Average pooling, 128     Meta-training LR 0.0001
@@
Model performance is measured as the average over 10,000 episodes sampled from the meta-test set. As is common in GFSL [59], we evaluate seen and unseen classes separately and report the harmonic mean of their scores. However, we adopt AUC–ROC instead of accuracy [54, 36] to align with evaluation standards in the medical domain.
We employ three metrics commonly used in GFSL [59], each reported with a 95% confidence interval: one computed for the seen classes, another for the unseen classes, and a third one for the harmonic mean ($HM$), which are defined as follows.
  • 1.
@@
    Note that the harmonic mean is commonly used in GFSL because it mitigates the dominance of seen classes in the overall performance [59].

### v2

@@ 1 Introduction
In recent decades, deep learning has revolutionized medical image analysis, particularly in the field of radiology [tfmedical, chextransfer2021, cherti2022]. Deep neural networks have enabled the processing of large volumes of radiological data, the extraction of complex features, and the development of models that can enhance the accuracy of medical diagnoses. Despite these advances, a major challenge arises when only limited annotated data are available, as deep learning models typically require large amounts of labeled data to achieve strong performance. This issue is especially relevant to tasks such as pathology classification in chest X-rays, where labeled data can be scarce and difficult to obtain. To address this limitation, early research has explored the use of the standard few-shot classification (SFSC) paradigm, which aims to train models capable of generalizing to new classes using only a few labeled examples per class. However, this paradigm differs significantly from the way pathologies manifest in practice. From a clinical perspective, the objective is not merely to classify entirely new disease categories, but rather to distinguish between a combination of known pathologies and previously unseen ones. This highlights the need for approaches that are more closely aligned with the complexities of clinical settings.
@@ 2.1 Deep learning for chest X-ray classification
Deep neural networks, coupled with large-scale datasets, have enabled significant progress in several computer vision fields. Over the past few years, in an effort to take advantage of deep neural networks, datasets of increasingly specialized domains have been made publicly available. For instance, in the medical domain, multiple chest X-ray datasets have been introduced, such as CheXpert [CheXpert2019], Chest X-ray8 [chestX-ray82017], Chest X-ray14 [chestX-ray82017], MIMIC [MIMIC-CXR2019], MIMIC-CXR-JPG [mimicjpg], OpenI [openI], and PadChest [PadChest]. These datasets have been fundamental in the development of deep learning models for chest X-ray analysis and generation tasks. In contrast to ImageNet [ImageNet], the scale of these datasets is at least one order of magnitude smaller. In addition, the distributions of these datasets are greatly heterogeneous, e.g. the number and kind of pathologies, the class imbalance, the collection and labeling procedure, the quality of the images, and the patient population.
With the introduction of these datasets, several studies addressing pathology classification from chest X-rays using deep learning started to appear. For binary classification (i.e. presence or absence), tuberculosiscls focused on tuberculosis identification, whereas pneumoniadetection targeted pneumonia identification. Since the X-ray of a patient may exhibit signs of multiple diseases, the identification of pathologies from chest X-rays has often been formulated as a multi-label classification problem. For instance, comparisonmodels2019 used a ResNet-50 [resnet] architecture to classify fourteen pathologies in the ChestX-ray14 dataset, where each X-ray could be assigned to more than one pathology. Similarly, CheXpert2019 compared various ConvNet architectures for multi-label chest X-ray classification using CheXpert, finding that DenseNet121 outperformed ResNet152 [resnet], Inception-v4 [inceptionv4], and SE-ResNeXt101 [se-resnex].
@@ 2.2 Standard transfer learning
Transfer learning (TL) is a cornerstone of deep learning for image analysis, since it can reduce the amount of data and the computational resources required to train a model for a target task by leveraging the representations learned from one or multiple source tasks. In TL, multiple strategies to adapt the representations from the source task to the target task have been proposed. In practice, the most widely used transfer strategy has been standard transfer learning, which consists of pre-training models using a conventional batch-based training (as opposed to other training schemes). Specifically, ImageNet pre-training has been a standard practice for a wide variety of natural image tasks, including classification [efficientnet, bit2020], segmentation [r-cnn, yolo] and object detection [fcn].
@@
Due to the widespread use of ImageNet in practice, multiple works have studied the transferability of the learned representations from ImageNet to other natural image tasks [featuresoff-the-shelf, Yosinski2014, kornblith2019better, zhou2021convnets]. Surprisingly, even though some studies have suggested that the source and target datasets must be closely related for an effective knowledge transfer [Yosinski2014], ImageNet pre-training has been used with seeming success for wildly dissimilar image domains (e.g. medical images [chestX-ray82017, transfusion2019, chextransfer2021]). In contrast, there are transferability studies in specific domains where ImageNet pre-training has not provided any improvement over random initialization [transfusion2019, tfmedical, cherti2022].
@@
Other transferability studies have focused on analyzing the effect of the architecture size and the scale of the training dataset on the effectiveness of STL. In intra-domain scenarios, where the source and target datasets are closely related, studies have been mainly focused on natural image datasets. For example, bit2020 and scalinvisiontsf analyzed how the pre-training dataset size and the architecture depth influence knowledge transfer when both the source and target datasets are composed of natural images. The results from these studies have consistently shown better performance with larger architectures and pre-training datasets.
@@
In inter-domain scenarios, where the source and target datasets belong to different domains (e.g., natural images and chest X-rays), existing studies are scarce, not very systematic and report mixed results. transfusion2019 did not find significant differences on chest X-ray and retinal image classification performance using a ResNet-50 architecture when comparing ImageNet-1k pre-training with random initialization. chextransfer2021 studied the effect of ImageNet-1k pre-training on chest X-ray classification performance using ConvNet architectures of different sizes. Their results showed a slight performance improvement when using deeper pre-trained architectures. On the other hand, tfmedical studied the influence of ImageNet-1k, ImageNet-21k, and JFT-300M pre-training on classification performance using ResNets of different sizes. The target tasks considered in this study were cancer identification from mammograms, pathology classification from chest X-rays, and skin condition from dermatological images. The results were far from conclusive, observing performance improvements with larger pre-training datasets and architectures only in some target tasks. Similarly, cherti2022 carried out a comparative study of ResNet models pre-trained on ImageNet-1k, ImageNet-21k, and a combination of different chest X-ray datasets for pathology classification. They reported small improvements in performance when models were pre-trained on larger source datasets and transferred to larger target datasets. However, no performance improvement was observed when transferring for smaller target datasets, regardless of the size of the pre-training dataset.
@@ 2.3 Few-shot classification
Meta-learning is a transfer learning strategy which aims to generate models that can be quickly adapted to new tasks [learntolearn]. As opposed to STL, in meta-learning, new tasks are commonly known as episodes and are typically small with respect to both the number of classes and the number of examples per class. The most widely studied problem in meta-learning is few-shot classification, which is a multi-class classification problem where a few examples per class are available for training (typically, 1 or 5).
The earliest works in meta-learning proposed methods for multi-class classification on natural image datasets [MAML, matching, protonet, relationet, setfeat], such as MiniImageNet [matching] and FC100 [meta_transfer]. These datasets are reduced versions of ImageNet and CIFAR100 and were created to facilitate episodic training.
More recent works have applied meta-learning to domain-specific problems, particularly through datasets of various medical imaging modalities. For instance, meta-learning methods have been studied in skin disease classification from dermatological images [metaderm2020], COVID-19 classification from chest CT scans [CHEN2021107826], and cancer classification from histological images [histopathological2019, shakeri2022fhist]. Moreover, meta-learning methods have also been used for image segmentation in CT scans, magnetic resonance images [Tang2021RecurrentMR], and dermatological images [KHADKA2022105227].
@@ 3 MetaChest dataset
Clinical data collection is a complex process involving several tasks which can require a considerable amount of time and resources. Data labeling is one of the tasks that can generate greater variability among chest X-ray datasets; the rightmost column in Table 1 summarizes the labeling strategy employed by each dataset. Most datasets derived annotations automatically from radiology reports using natural language processing (NLP) methods, with the exception of PadChest which was annotated by expert radiologists. The specific strategy and tool used for annotating the chest X-rays directly influences the distribution of labels. For instance, MIMIC provides two sets of labels that have different distributions: one generated by NegBio [negbio] and another by CheXpert [CheXpert2019].
@@
Table 1: Comparison of publicly available chest X-ray datasets. Here, MIMIC [MIMIC-CXR2019] refers to MIMIC-CXR-JPG [mimicjpg].
@@ 3.1 Data
The frequency of each pathology in MetaChest is shown in Figure 1. As can be observed, there is a pronounced class imbalance, with the most frequent pathology (Effusion) occurring nearly two orders of magnitude more often than the least frequent one (Hernia). With respect to labeling, MetaChest has a label cardinality (average number of labels per image) of 1.84 and a label density (label over the total number of labels; see Tsoumakas2010) of 0.12.
@@ 3.2 Meta-learning partition
Although cherti2022 used a chest X-ray dataset that combined multiple datasets, it is not publicly available, its generation procedure is not described, and it does not provide appropriate partitions for meta-learning evaluation. Similarly, TorchXRayVision [Cohen2022xrv] is a library that allows the combination of different chest X-ray datasets, but it does not consider SFSL scenarios either. Conversely, MetaChest employs a disjoint class partition, enabling experimentation in SFSL settings. In addition, $\mathcal{C}_{meta-tst}$ is composed of the classes with the fewest examples available across the four original datasets, which is convenient for evaluating classification methods on images collected from multiple hospitals.
@@ 4.1 Standard transfer learning
Standard transfer learning (STL) is the most widely studied and spread strategy for computer vision tasks. When performing STL, we can identify two main stages [tfsurvey]:
@@ 4.2 Meta-learning
Meta-learning (MTL) is an alternative paradigm to STL, which aims to obtain models that can adapt to novel tasks with unseen classes and very few examples per class [tfsurvey]. In other words, the objective of MTL is to achieve a more efficient transfer in terms of data. Similar to STL, the transfer process of MTL consists of a pre-training stage followed by an adaptation stage. Although in MTL these stages are commonly referred to as meta-training and meta-test [optLSTM], for the sake of consistency and clarity, we use the terms pre-training and adaptation for both STL and MTL. In this work, we are interested in two MTL formulations for classification: standard few-shot classification (SFSC) and generalized few-shot classification.
@@ 5.2.1 ProtoNet-ML
ProtoNet [protonet] is a multi-class classification method that has been widely studied in the SFSC literature. In this work, we propose an extension to handle multi-label classification, which we call ProtoNet-ML. Following the original method, ProtoNet-ML computes a $D$-dimensional prototype $\mathbf{z}_{c}\in\mathbb{R}^{D}$ for each class $c$ as follows:
@@
Beyond the Euclidean distance, ProtoNet-ML can be instantiated with other functions, including the Minkowski distance and, with slight modifications, the cosine distance. Moreover, ProtoNet-ML is a flexible method that supports arbitrary activation functions and can operate directly on logits. In our experiments, however, we employ the sigmoid function, as it is the conventional and most natural choice for binary classification.
Figure 4: Representation subspaces for multi-class ProtoNet (left) and ProtoNet-ML (right). In multi-class ProtoNet, a representation vector $g_{\phi}(\mathbf{x})$ is associated with only one prototype, whereas in ProtoNet-ML, a representation vector $f_{\phi}(\mathbf{x})$ may be associated with one or more prototypes. Note that $f_{\phi}(\cdot)$ is the network backbone followed by an encoding layer to reduce representation vector dimensionality.
@@ 5.2.2 BatchBased
BatchBased is a method inspired by [chen2019closer] that employs STL-based training while maintaining MTL-based episode evaluation. On top of the backbone $f_{\phi}$, BatchBased adds a head module $g_{\varphi}$ (single fully connected layer) with trainable parameters $\varphi$. The probability distribution for a test example $\mathbf{x}$ given a class $c$ is computed as:
@@ 6 Results and discussion
In this section, we analyze the adaptation process of different models across various formulations of the multi-label classification task using few examples per pathology. First, we compare how two distinct learning paradigms leverage ImageNet pre-training. Next, we examine their behavior across a broad set of few-shot learning tasks designed to reflect challenges commonly encountered in medical settings. In addition, we investigate factors that influence the adaptation process, including image resolution, and variations in architectural connectivity patterns. Finally, we study the influence of hyperparameters on classification performance and provide illustrative examples of the resulting model predictions.
@@ 6.1 Experimental setup
Parameter Configuration Data     Distribution Complete     Image size 384 Task     Training batch size 64     Training $n$-way, $k_{trn}$ $k_{tst}$ 3, 30, 30     Validation $n$-way, $k_{trn}$ $k_{tst}$ 3, 30, 30     Test $n$-way, $k_{trn}$ $k_{tst}$ 3, 30, 30 Backbone     Architecture MobileNetV3Small075     Pre-training I1K Training     Meta-trn, meta-val, meta-tst episodes 1,000, 100, 10,000     Max epochs 150     Optimizer AdamW     Stop metric, patience $HM$, 10     Float precision 16bit BatchBased     Meta-trn LR 0.0001     Meta-val $t_{steps}$, $ptc_{trn}$, $lr_{head}$ 100, 0.5, 0.05     Meta-tst $t_{steps}$, $ptc_{trn}$, $lr_{head}$ 100, 0.5, 0.05 ProtoNet-ML     Encoding layer type, size Average pooling, 128     Meta-training LR 0.0001
@@
Model performance is measured as the average over 10,000 episodes sampled from the meta-test set. As is common in GFSL [gsfl], we evaluate seen and unseen classes separately and report the harmonic mean of their scores. However, we adopt AUC–ROC instead of accuracy [chestclassi, ClassificationCausal] to align with evaluation standards in the medical domain.
We employ three metrics commonly used in GFSL [gsfl], each reported with a 95% confidence interval: one computed for the seen classes, another for the unseen classes, and a third one for the harmonic mean ($HM$), which are defined as follows.
@@
    Note that the harmonic mean is commonly used in GFSL because it mitigates the dominance of seen classes in the overall performance [gsfl].